A stock is expected to pay a dividend of $1.00 at the end of the year (i.e., D1 = $1.00), and it should continue to grow at a constant rate of 5% a year. If its required return is 13%, what is the stock's expected price 1 year from today? Do not round intermediate calculations. Round your answer to the nearest cent.

$  

___________________________________________________________________________________--

Holtzman Clothiers's stock currently sells for $39.00 a share. It just paid a dividend of $1.00 a share (i.e., D0 = $1.00). The dividend is expected to grow at a constant rate of 7% a year.

What stock price is expected 1 year from now? Round your answer to the nearest cent.
$  

What is the required rate of return? Do not round intermediate calculations. Round your answer to two decimal places.
  %
Answer #1

a.Current price=D1/(Required return-Growth rate)

=1/(0.13-0.05)

=$12.5

Hence expected price=Current price*(1+Growth Rate)

=12.5*1.05

=$13.13(Approx).

b.Expected price=Current price*(1+Growth Rate)

=39*1.07

=$41.73

Required rate=(D1/Current price)+Growth rate

=[(1*1.07)/39]+0.07

=9.74%(Approx).